Give an example of a straight line passing through point (1,3,4) and parallel to the plane X+Y+Z=1

Grade:

1 Answers

vikas askiitian expert
13 years ago

there are infinite examples for this question

L1 : x-1/1 = y-3/1 = z-4/-2 

L2 :  x-1/-2 = y-3/1= z-4/1

L3 : x-1/1 = y-3/-2 = z-4/1

L4 : x-1/2  = y-3/3 = z-4/-5

  these all are possible examples